# A deep learning model for chemical shieldings in molecular organic solids including anisotropy

### v1

@@ Abstract
Nuclear Magnetic Resonance (NMR) chemical shifts are powerful probes of local atomic and electronic structure that can be used to resolve the structures of powdered or amorphous molecular solids. Chemical shift driven structure elucidation depends critically on accurate and fast predictions of chemical shieldings, and machine learning (ML) models for shielding predictions are increasingly used as scalable and efficient surrogates for demanding ab initio calculations. However, the prediction accuracies of current ML models still lag behind those of the DFT reference methods they approximate, especially for nuclei such as 13C and 15N. Here, we introduce ShiftML3, a deep-learning model that improves the accuracy of predictions of isotropic chemical shieldings in molecular solids, and does so while also predicting the full shielding tensor. On experimental benchmark sets, we find root-mean-squared errors with respect to experiment for ShiftML3 that approach those of DFT reference calculations, with RMSEs of 0.53 ppm for 1H, 2.4 ppm for 13C, and 7.2 ppm for 15N, compared to DFT values of 0.49 ppm, 2.3 ppm, and 5.8 ppm, respectively.
@@
In a nutshell, NMR crystallography consists in experimentally measuring and assigning chemical shifts for each of the atomic sites in a structure, and then comparing the measured shifts to those predicted by computation for a pool of candidate structures, that are themselves enumerated by a computational search. Chemical shifts are traditionally calculated using DFT methods. [7, 11, 12, 13, 14] This workflow has been successful for materials ranging from hybrid-perovskites [15, 16] and inorganic materials [17, 18, 19, 20, 21, 22] to enzyme activate sites, [23, 24, 10] organic molecular solids, [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and notably for amorphous drug formulations. [40, 41, 42, 43] One of the key bottlenecks to address in this workflow is both the computational cost and the accuracy of chemical shift predictions.
@@
Here, we present a new machine learning model (dubbed ShiftML3) that predicts the full chemical shielding tensors, trained using shieldings calculated at the GIPAW/PBE level of theory. The model is based on an ensemble of Point Edge Transformer (PET) models. [61] The model not only predicts anisotropic shieldings, but also yields significantly improved accuracy with respect to DFT, with up to a two-fold improvement over ShiftML2 for the isotropic shieldings of heavy atoms.
@@
Here we use the training set from the most recent previous ShiftML model, and for which the computed shifts (computed at the PBE level of theory) contain the full tensor information. [49] This dataset is composed of molecular solids where the structures include (up to) 12 elements [H, C, N, O, F, Na, Mg, P, S, Cl, K, Ca] commonly found in organic compounds. It contains a total of 14,146 structures, including 7,548 sampled from the Cambridge structural database (CSD) via farthest point sampling (FPS) [70] to cover a broad range of structural motifs, and 6,598 additional thermally distorted structures. The ShiftML2 test set was divided into a new validation (505 structures, with 41,599 atomic environments) and a new test set (1,254 structures, with 106,230 atomic environments). In this split we choose a stratified selection strategy that ensures that both the distorted and the relaxed structure of any one CSD identifier are found either in the validation or the test set, avoiding data leakage between the sets. Furthermore, we ensure that both the new validation and new test sets contain at least a few environments of all 12 chemical elements. To compare the accuracies between ShiftML2 and ShiftML3 in a consistent fashion, the accuracies are reported on this new test set (referred to as the CSD-test set). More details of the training, validation and test sets for each element in CSD-test are given in section S1 of the supporting information. Due to computational constraints, ShiftML2 was trained on a representative subselection of atomic environments from the training set. The favorable computational scaling of neural network training via stochastic gradient descent allowed us to train ShiftML3 using all 1.4 million atomic environments contained in the 14,146 structures in the training set. The updated dataset containing the chemical shielding tensor data is available via a public data record associated with this publication. [71] We also use established datasets of experimental isotropical chemical shielding values, with 13 structures for 1H, 21 for 13C, and 15 for 15N, [49, 14] and a benchmark set of experimental 13C CSA principal values for 19 compounds. [72]
@@
Figure 1: Parity plots between DFT-computed shieldings and predictions using ShiftML2 or ShiftML3 for (a,b) 1H, (c,d) 13C and (e,f) 15N evaluated on the new CSD-test set. The black line shows perfect correlation while the colour each point is correlated to the prediction uncertainty of that atomic site. Each point is colored according to the rescaled prediction uncertainty associated in each atomic site.
To evaluate the performance of ShiftML3, we first compare predicted isotropic chemical shieldings against DFT calculated isotropic shieldings on the CSD-test set. Figure 1 shows parity plots illustrating the overall agreement between ShiftML3 and DFT with an RMSE of 0.43 ppm, 2.3 ppm and 10.4 ppm for 1H, 13C and 15N, respectively. These correspond to an error reduction between 20% to a factor of 2 relative to ShiftML2. Similar levels of improvement are seen also for almost all nuclei represented in the dataset, as shown in Table 1 (with the exception of 39K, for which there are only 6 atomic enviroments in the test set, making the assessment inconclusive).
@@
Figure 2: Parity plots between experimental chemical shifts and predictions from ShiftML2, ShiftML3 and DFT calculations for (a,b,c) 1H, (d, e, f) 13C, and (g, h, i)15N evaluated on the benchmark structures from Ramos et al. [14] for 13C and 15N, and the 13 structures where ShiftML2 was evaluated for 1H. The black line shows the linear regression between experimental chemical shifts and predicted isotropic shiedings with their slope and intercept indicated in each respective subplot.
@@
The improvements in prediction accuracy against the DFT reference also translate to an increased accuracy of ShiftML predictions against experiments. We find that on a representative benchmark set of experimentally measured 1H and 13C shieldings, [49, 14] ShiftML3 predicts chemical shieldings with accuracies almost on par with GIPAW DFT, namely: 0.53(0.49) ppm for 1H and 2.4(2.3) ppm for 13C, for ShiftML3(DFT). Figure 2, shows parity plots of isotropic chemical shieldings estimated with ShiftML2, ShiftML3 and DFT with respect to experimental chemical shifts. DFT values were evaluated recomputing the NMR shieldings using GIPAW [7] on the the Quantum ESPRESSO software [90, 91] with the same DFT parameters as in Ref. 49. The conversion between experimental chemical shifts and predicted shieldings is performed by fitting globally the shieldings across all test compounds to a linear form, also shown in figure 2. Further details about the experimental benchmark are discussed in section S7 of the supporting information.
@@ Acknowledgements.
The supporting information is provided below, containing additional details on dataset and model construction, uncertainty quantification, and the usage of the ShiftML3 Python package, which is available under an open source license ad https://github.com/lab-cosmo/shiftml. The training, validation and testing datasets containing chemical shielding tensor data as well as training input files, model files and evaluation scripts will be made available upon publication via the public data record associated with this publication. [71]
@@
We benchmark the evaluation time of the ShiftML3 model. All benchmark timings are obtained via the ASE interface of the ShiftML library. We report timings that include setting up the ASE calculator once and then computing chemical shielding tensors for a series of structures. We believe that this protocol reflects most accurately the evaluation timings for a common NMR crystallography workflow. The timings via the ASE interface include also the neighbour list construction. The ASE interface does not support batched evaluations, or efficient parallelization over ensemble models for smaller system sizes. We report CPU-only and CPU+GPU timings. We report evaluation times on a Lenovo workstation with a AMD Ryzen Threadripper PRO 3955WX CPU with 16 cores. CPU+GPU timings are obtained on the same workstation, using a NVIDIA GeForce RTX 4070 Ti-super consumer GPU, which is representative of a configuration that does not require access to high-end AI-dedicated hardware. As shown in Table S3, even a low-end GPU accelerates by 4x-6x the evaluation of structures with the size of typical small-molecule pharmaceuticals. The acceleration is also evident at larger system sizes: the scaling of the ShiftML3 model for diamond supercells of increasing size is almost perfectly linear, and there is an order-of-magnitude acceleration when using a GPU (Fig. S2). These timings are also representative of large biomolecules, as hinted at by the inference timings for the protein crystal of 1PGA [92], the B1 immunoflobulin-binding domain of the streptococcal protein G with 56 amino acid residues and 3,656 atoms in the conventional cell (also shown in Fig. S2). 1PGA was a target of several solid-state NMR studies [93, 94] and therefore lies well within the range of atomistic systems of interest for which chemical shielding predictions may complement experimental studies. A computation of chemical shieldings for this protein takes 5.6 s using the GPU. Moving to even larger systems, we find that on the consumer GPU used for benchmarking with 16GB of VRAM, we can predict 13C shieldings in diamond supercells of up to 16,000 atoms before exhausting the graphics card video memory, demonstrating the memory efficiency of ShiftML3.
@@
Set Time per struct. [s] CPU/GPU $\overline{\text{N}}_{\text{atoms}}$
@@
The ShiftML3 model integrates into existing materials modelling and crystal structure prediction workflows via the atomic simulation environment. Chemical shielding predictions can be obtained by initializing an ASE calculator object and reading in the geometry file in one of the various supported file formats. Once the calculator is initialized, the model weights are automatically downloaded and loaded by the model. A minimalistic usage example is given below, and a more thorough tutorial is available at http://atomistic-cookbook.org/shiftml/shiftml-example.html.
@@
The final, calibrated, uncertainties are then obtained as $\alpha^{2}\sigma^{2}_{\text{pred}}(A)$ although in practice we scale the ensemble members around their mean, which simplifies uncertainty propagation. [86] We find that for isotropic chemical shieldings and shielding tensor principal components the calibration factors $\alpha$ are moderate and always close to $1$, indicating that the ShiftML3 ensemble uncertainty estimates, are already well calibrated.
@@
In Figure. S6 we show parity plots of symmetric shielding tensor component predictions and principal tensor component predictions.
@@
Evaluation of 1H was done on the same 13 structures where ShiftML2 was evaluated [49] on the same geometry optimized structures. Predicted shieldings of methyl or primary amine groups were averaged out accounting for rapid rotation, resulting in the averaged sites have equivalent chemical shielding. Ambiguous experimental assignments were solved by iterating along all possible permutations of those unassigned shifts and taking the assignment with the lowest RMSE between predicted and experimental chemical shift such as for methylene groups. Only the permutation with the lowest RMSE entered the global linear regression. We note that the reported RMSE comparing prediction to experiment for ShiftML2 is modestly different than the original work. This difference originates from the use of a global regression as compared to on-the-fly regression used in the original work. For 13C and 15N, the accuracy was were evaluated from the benchmark consisting of of 36 structures from Ramos et al. [14] on geometries optimized on the PBE-D3(BJ)level of theory [97]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72]

### v2

@@ Abstract
Nuclear Magnetic Resonance (NMR) chemical shifts are powerful probes of local atomic and electronic structure that can be used to resolve the structures of powdered or amorphous molecular solids. Chemical shift driven structure elucidation depends critically on accurate and fast predictions of chemical shieldings, and machine learning (ML) models for shielding predictions are increasingly used as scalable and efficient surrogates for demanding ab initio calculations. However, the prediction accuracies of current ML models still lag behind those of the DFT reference methods they approximate. Here, we introduce ShiftML3, a deep-learning model that improves the accuracy of predictions of isotropic chemical shieldings in molecular solids, and does so while also predicting the full shielding tensor. On experimental benchmark sets, we find root-mean-squared errors with respect to experiment for ShiftML3 that approach those of DFT reference calculations, with RMSEs of 0.53 ppm for 1H, 2.4 ppm for 13C, and 7.2 ppm for 15N, compared to DFT values of 0.49 ppm, 2.3 ppm, and 5.8 ppm, respectively.
@@
In a nutshell, NMR crystallography consists in experimentally measuring and assigning chemical shifts for each of the atomic sites in a structure, and then comparing the measured shifts to those predicted by computation for a pool of candidate structures, that are themselves enumerated by a computational search. Chemical shifts are traditionally calculated using DFT methods. [7, 11, 12, 13, 14] This workflow has been successful for materials ranging from hybrid-perovskites [15, 16] and inorganic materials [17, 18, 19, 20, 21, 22] to enzyme active sites, [23, 24, 10] organic molecular solids, [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and notably for amorphous drug formulations. [40, 41, 42, 43] One of the key bottlenecks to address in this workflow is both the computational cost and the accuracy of chemical shift predictions.
@@
Here, we present a new machine learning model (dubbed ShiftML3) that predicts the full chemical shielding tensors, trained using shieldings calculated at the GIPAW/PBE level of theory. The model is based on an ensemble of Point Edge Transformer (PET) models. [61] The model not only predicts anisotropic shieldings, but also yields significantly improved accuracy with respect to DFT, with up to a two-fold improvement over ShiftML2 for the isotropic shieldings of nuclei with Z >1.
@@
Here we use the training set from the most recent previous ShiftML model, and for which the computed shifts (computed at the PBE level of theory) contain the full tensor information. [49] This dataset is composed of molecular solids where the structures include (up to) 12 elements [H, C, N, O, F, Na, Mg, P, S, Cl, K, Ca] commonly found in organic compounds. It contains a total of 14,146 structures, including 7,548 sampled from the Cambridge structural database (CSD) via farthest point sampling (FPS) [70] to cover a broad range of structural motifs, and 6,598 additional thermally distorted structures. The ShiftML2 test set was divided into a new validation (505 structures, with 41,599 atomic environments) and a new test set (1,254 structures, with 106,230 atomic environments). In this split we chose a stratified selection strategy that ensures that both the distorted and the relaxed structure of any one CSD identifier are found either in the validation or the test set, avoiding data leakage between the sets. Furthermore, we ensure that both the new validation and new test sets contain at least a few environments of all 12 chemical elements. To compare the accuracies between ShiftML2 and ShiftML3 in a consistent fashion, the accuracies are reported on this new test set (referred to as the CSD-test set). More details of the training, validation and test sets for each element in CSD-test are given in section S1 of the supporting information. Due to computational constraints, ShiftML2 was trained on a representative subselection of atomic environments from the training set. The favorable computational scaling of neural network training via stochastic gradient descent allowed us to train ShiftML3 using all 1.4 million atomic environments contained in the 14,146 structures in the training set. The updated dataset containing the chemical shielding tensor data is available via a public data record associated with this publication. [71] We also use established datasets of experimental isotropic chemical shielding values, with 13 structures for 1H, 21 for 13C, and 15 for 15N, [49, 14] and a benchmark set of experimental 13C CSA principal values for 19 compounds. [72]
@@
Figure 1: Parity plots between DFT-computed shieldings and predictions using ShiftML2 or ShiftML3 for (a,b) 1H, (c,d) 13C and (e,f) 15N evaluated on the new CSD-test set. The black line shows perfect correlation. Each point is colored according to the rescaled prediction uncertainty associated in each atomic site.
To evaluate the performance of ShiftML3, we first compare predicted isotropic chemical shieldings against DFT calculated isotropic shieldings on the CSD-test set. Figure 1 shows parity plots illustrating the overall agreement between ShiftML3 and DFT for the nuclei that are most often used in NMR crystallography. We observe an RMSE of 0.43 ppm, 2.3 ppm and 10.4 ppm for 1H, 13C and 15N, respectively. These correspond to an error reduction relative to ShiftML2 between 20% for 1H, to a factor of 2 for 13C. Similar levels of improvement are seen also for almost all nuclei represented in the dataset, as shown in Table 1 (with the exception of 39K, for which there are only 6 atomic enviroments in the test set, making the assessment inconclusive).
@@
Figure 2: Parity plots between experimental chemical shifts and predictions from ShiftML2, ShiftML3 and DFT calculations for (a,b,c) 1H, (d, e, f) 13C, and (g, h, i)15N evaluated on the benchmark structures from Ramos et al. [14] for 13C and 15N, and the 13 structures where ShiftML2 was evaluated for 1H. The black line shows the linear regression between experimental chemical shifts and predicted isotropic shieldings with their slope and intercept indicated in each respective subplot.
@@
The improvements in prediction accuracy against the DFT reference also translate to an increased accuracy of ShiftML predictions against experiments. We find that on a representative benchmark set of experimentally measured 1H and 13C shieldings, [49, 14] ShiftML3 predicts chemical shieldings with accuracies almost on par with GIPAW DFT, namely: 0.53(0.49) ppm for 1H and 2.4(2.3) ppm for 13C, for ShiftML3(DFT). Given that the errors of ShiftML3 against DFT are of similar size, this hints at substantial error cancellation. We also note that both computational predictions are affected by non-negligible error contributions due to the use of a static structure, neglecting finite-temperature effects and quantum fluctuations [90]. Figure 2, shows parity plots of isotropic chemical shieldings estimated with ShiftML2, ShiftML3 and DFT with respect to experimental chemical shifts. DFT values were evaluated recomputing the NMR shieldings using GIPAW [7] on the the Quantum ESPRESSO software [91, 92] with the same DFT parameters as in Ref. 49. Quantum Espresso GIPAW calculations for these 48 structures took approximately 80 hours on 32 cores of a modern compute node for all three benchmark sets. In contrast, ShiftML3 yielded shielding predictions for all structures in just 41 seconds on a personal computer (MacBook Pro M1), demonstrating speed-ups of over three orders of magnitude. The conversion between experimental chemical shifts and predicted shieldings is performed by fitting globally the shieldings across all test compounds to a linear form, also shown in figure 2. Further details about the experimental benchmark are discussed in section S7 of the supporting information.
@@ Acknowledgements.
The supporting information is provided below, containing additional details on dataset and model construction, uncertainty quantification, and the usage of the ShiftML3 Python package, which is available under an open source license at https://github.com/lab-cosmo/shiftml. The training, validation and testing datasets containing chemical shielding tensor data as well as training input files, model files and evaluation scripts will be made available upon publication via the public data record associated with this publication. [71]
@@
We benchmark the evaluation time of the ShiftML3 model. All benchmark timings are obtained via the ASE interface of the ShiftML library. We report timings that include setting up the ASE calculator once and then computing chemical shielding tensors for a series of structures. We believe that this protocol reflects most accurately the evaluation timings for a common NMR crystallography workflow. The timings via the ASE interface include also the neighbour list construction. The ASE interface does not support batched evaluations, or efficient parallelization over ensemble models for smaller system sizes. We report CPU-only and CPU+GPU timings. We report evaluation times on a Lenovo workstation with a AMD Ryzen Threadripper PRO 3955WX CPU with 16 cores. CPU+GPU timings are obtained on the same workstation, using a NVIDIA GeForce RTX 4070 Ti-super consumer GPU, which is representative of a configuration that does not require access to high-end AI-dedicated hardware. As shown in Table S3, even a low-end GPU accelerates by 4x-6x the evaluation of structures with the size of typical small-molecule pharmaceuticals. The acceleration is also evident at larger system sizes: the scaling of the ShiftML3 model for diamond supercells of increasing size is almost perfectly linear, and there is an order-of-magnitude acceleration when using a GPU (Fig. S2). These timings are also representative of large biomolecules, as hinted at by the inference timings for the protein crystal of 1PGA [93], the B1 immunoglobulin-binding domain of the streptococcal protein G with 56 amino acid residues and 3,656 atoms in the conventional cell (also shown in Fig. S2). 1PGA was a target of several solid-state NMR studies [94, 95] and therefore lies well within the range of atomistic systems of interest for which chemical shielding predictions may complement experimental studies. A computation of chemical shieldings for this protein takes 5.6 s using the GPU. Moving to even larger systems, we find that on the consumer GPU used for benchmarking with 16GB of VRAM, we can predict 13C shieldings in diamond supercells of up to 16,000 atoms before exhausting the graphics card video memory, demonstrating the memory efficiency of ShiftML3.
@@
Set
Time per struct. [s]
CPU/GPU
$\overline{\text{N}}_{\text{atoms}}$
@@
The ShiftML3 model integrates into existing materials modelling and crystal structure prediction workflows via the Atomic Simulation Environment (ASE). Chemical shielding predictions can be obtained by initializing an ASE calculator object and reading in the geometry file in one of the various supported file formats. Once the calculator is initialized, the model weights are automatically downloaded and loaded by the model. A minimalistic usage example is given below, and a more thorough tutorial is available at http://atomistic-cookbook.org/shiftml/shiftml-example.html.
@@
The final, calibrated, uncertainties are then obtained as $\alpha^{2}\sigma^{2}_{\text{pred}}(A)$ although in practice we scale the ensemble members around their mean, which simplifies uncertainty propagation. [86] We find that for isotropic chemical shieldings and shielding tensor principal components the calibration factors $\alpha$ are moderate and always close to 111, indicating that the ShiftML3 ensemble uncertainty estimates, are already well calibrated.
@@
In Figure S6 we show parity plots of symmetric shielding tensor component predictions and principal tensor component predictions.
@@
In Figure S7 we show parity plots of the prinicpal tensor components resolved for individual principal components $\sigma_{11}$,$\sigma_{22}$ and $\sigma_{33}$. In Table S5 we list prediction accuracies of the individual principal components.
Figure S7: Component resolved parity plots of the principal tensor components of 15N shielding tensors. Combined principal components are plotted in a), $\sigma_{\text{11}}$ in b), $\sigma_{\text{22}}$ in c) and $\sigma_{\text{33}}$ in d).
Table S5: Predicted accuracy of $\sigma_{\text{PAS}}$ of (sML3) on the CSD-test hold-out set, resolved for individual principal components $\sigma_{\text{11}}$, $\sigma_{\text{22}}$ and $\sigma_{\text{33}}$.
Species MAE [ppm] RMSE [ppm] RMSE [%]
$\sigma_{11}$ $\sigma_{22}$ $\sigma_{33}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{33}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{33}$
${}^{1}\text{H}$ 0.65 0.60 0.69 0.86 0.77 0.90 21 22 21
${}^{13}\text{C}$ 2.77 2.96 2.19 4.17 4.28 3.07 5 8 10
${}^{15}\text{N}$ 9.55 6.42 5.66 21.10 11.32 12.57 11 9 19
${}^{17}\text{O}$ 14.40 9.35 7.09 23.70 13.28 10.41 8 6 13
${}^{19}\text{F}$ 6.02 6.35 4.83 8.99 9.40 7.45 11 10 20
${}^{33}\text{S}$ 26.75 19.35 21.50 57.02 32.32 34.40 29 22 21
${}^{31}\text{P}$ 12.87 11.43 15.75 23.90 17.49 28.15 28 25 45
${}^{35}\text{Cl}$ 14.75 14.00 13.31 20.99 21.39 19.03 11 14 32
${}^{23}\text{Na}$ 4.76 5.27 2.82 4.95 6.79 3.40 68 65 43
${}^{43}\text{Ca}$ 4.66 4.98 3.96 4.74 5.48 4.52 23 30 34
${}^{25}\text{Mg}$ 3.68 4.97 10.17 4.97 5.39 12.45 20 23 45
${}^{39}\text{K}$ 5.22 3.31 4.74 6.12 3.78 5.43 79 42 50
@@
Evaluation of 1H was done on the same 13 structures on which ShiftML2 was evaluated [49], using the same geometry-optimized structures. Predicted shieldings of methyl or primary amine groups were averaged to account for rapid rotation, resulting in equivalent chemical shielding for the averaged sites. Ambiguous experimental assignments were resolved by iterating over all possible permutations of the unassigned shifts—such as for methylene groups—and selecting the assignment with the lowest RMSE between predicted and experimental chemical shifts. Only the permutation with the lowest RMSE entered the global linear regression. We note that the reported RMSE comparing prediction to experiment for ShiftML2 is modestly different from the original work. This difference originates from the use of a global regression compared with the on-the-fly regression used in the original work.
For 13C and 15N, the accuracies were evaluated using the benchmark of 36 structures from Ramos et al. [14], with geometries optimized at the PBE-D3(BJ) level of theory [98]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72].
Evaluation of 1H was done on the same 13 structures where ShiftML2 was evaluated [49] on the same geometry optimized structures. Predicted shieldings of methyl or primary amine groups were averaged out accounting for rapid rotation, resulting in the averaged sites have equivalent chemical shielding. Ambiguous experimental assignments were solved by iterating along all possible permutations of those unassigned shifts and taking the assignment with the lowest RMSE between predicted and experimental chemical shifts such as for methylene groups. Only the permutation with the lowest RMSE entered the global linear regression. We note that the reported RMSE comparing prediction to experiment for ShiftML2 is modestly different than the original work. This difference originates from the use of a global regression as compared to on-the-fly regression used in the original work. For 13C and 15N, the accuracy was were evaluated from the benchmark consisting of of 36 structures from Ramos et al. [14] on geometries optimized on the PBE-D3(BJ)level of theory [98]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72]